OpenStudy (anonymous):

Prove that x^2 + 1/x^2 >= 2

OpenStudy (dumbcow):

multiply by x^2 set equal to 0 \[x^{4} -2x^{2} +1 \ge 0\] Factor \[(x^{2} -1)^{2} \ge 0\] this is True , everything squared is greater than 0 thus original statement is True

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

or go other way round we know that anything squared is always >=0 A^2 > = 0 put A = x+1/x solve